OpenStudy (anonymous):

Solve each system by the addition method 3x+5y=21 -9x+4y=-6 put your answer in ordered pair (x,y) form

OpenStudy (anonymous):

3 (3x+5y=21) -9x+4y=-6 9x + 15y = 63 -9x + 4y = -6 19y = 57 y = 3 3x+5y=21 3x+5(3)=21 3x + 15 = 21 3x = 6 x = 2 (2, 3)